If B and C are the real roots of x^2+Bx+C=0 where B does NOT equal 0 and where C does not equal 0, find the product of B and C.
Answer by ptaylor(2198)

x^2+Bx+C=0
we are told that:
(x-B)(x-C)=0 expand using FOIL
x^2-Cx-Bx+BC=0 rewrite as follows:
x^2+(-C-B)x+BC=0
-C-B=B
-C=2B
C=-2B
and from the third term in the quadratic, we see that:
B=1
which means that C=-2 and B=1
therefore B*C=-2*1=-2
hope this helps---ptaylor